4 ms·
> a 1% chance per year, that implies a 63% probability per century. I'm not sure I'm following the math here?
by anhner 4y ago
> a 1% chance per year, that implies a 63% probability per century.
I'm not sure I'm following the math here?
- Synaesthesia 4y agoMaybe cause nuclear weapons only existed for 70 years ...?
- schiffern 4y agoWho said otherwise? Putting it in terms of "probability per century" is intended for forward-looking projection, not backwards.
- tralarpa 4y ago1 - (1 - 0.01)^100
- anhner 4y agoThanks, now I understand how they reached the number but not why it's so convoluted. I only took one stats course in college. So it's P(bomb exploding in a century) = 1 - P(bomb not exploding in a century) = 1 - (P(bomb not exploding in a year)) ^ 100 = 1 - (1 - P(bomb exploding in a year)) ^ 100 Why is it not something simpler, like P(bomb exploding in a year) ^ 100 ?
- tralarpa 4y ago> Why is it not something simpler, like P(bomb exploding in a year) ^ 100 ? That would be "bomb exploding EVERY year (= this year AND next year AND ...)". What we actually want is "bomb exploding this year OR bomb exploding next year OR bomb exploding this year and next year OR bomb exploding in two years OR ...". But that's too complicated, so we use the trick with 1 - P(bomb not exploding in a century).
- anhner 4y agoThat makes sense. Thanks for taking the time to clear up a stranger's confused mind! :)